OpenStudy (anonymous):

Lim of - (x-9)/root(x) -3 as x approaches 9

OpenStudy (yttrium):

You know conjugation?

OpenStudy (anonymous):

yes, i get 0? but i am not sure.

OpenStudy (anonymous):

will i have to do Lopitals?

random231 (random231):

yes as it is 0/0 form you can apply lhospitals rule.

OpenStudy (anonymous):

but wont i get 0/9 since i did the conjugate?

OpenStudy (anonymous):

nevermind, i got it why

OpenStudy (anonymous):

do i take the derivative of the original equation? or the conjugated one?

OpenStudy (yttrium):

Ooopsie. I see a better solution. Is it \[\lim_{x \rightarrow 9}-\frac{ (x-9) }{ \sqrt{x}-3 } \]

OpenStudy (anonymous):

yes that is the original

OpenStudy (yttrium):

Okay okay. Do you know that\[x-9 = (\sqrt{x} - 3)(\sqrt{x}+3)\]

OpenStudy (yttrium):

And with that we can cancel one term which is \[\sqrt{x}-3\]

OpenStudy (yttrium):

Hence our limit will then become \[\lim_{x \rightarrow 9} - (\sqrt{x}+3)\]

OpenStudy (yttrium):

Do you get my point?

OpenStudy (anonymous):

Yes, so we are allowed to do that?

OpenStudy (yttrium):

Of course :))

OpenStudy (anonymous):

iget -6?

OpenStudy (yttrium):

That's right.

OpenStudy (anonymous):

but it doesn't seem like it has a limit of -6?